TERAHERTZ SPECTROSCOPY AND GLOBAL ANALYSIS OF THE BENDING VIBRATIONS OF ^{12}C_{2}H_{2} and ^{12}C_{2}D_{2}

\underline{SHANSHAN YU}, BRIAN J. DROUIN, JOHN C. PEARSON AND HERBERT M. PICKETT, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109; VALERIO LAT-TANZI AND ADAM WALTERS, Centre d’Etude Spatiale des Rayonnements, Universit de Toulouse [UPS], CNRS [UMR 5187], 9 avenue du Colonel Roche, BP 44346, F-31028 Toulouse Cedex 4, France.

Symmetric molecules have no permanent dipole moment and are undetectable by rotational spectroscopy. Their interstellar observations have previously been limited to mid-infrared vibration-rotation spectroscopy. Although relatively weak, vibrational difference bands provide a means for detection of non polar molecules by terahertz techniques with microwave precision. Herschel, SOFIA, and ALMA have the potential to identify a number of difference bands of light symmetric species, e.g., C_{2}H_{2}, CH_{4} and C_{3}. This paper reports the results of the laboratory study on ^{12}C_{2}H_{2} and ^{12}C_{2}D_{2}. The symmetric isotopomers of acetylene have two bending modes, the trans bending ν_{4} (^{1}π_{g}), and the cis bending ν_{5} (^{1}π_{u}). For ^{12}C_{2}H_{2}, the two bending modes occur at 612 and 729 cm^{−1}, respectively. For ^{12}C_{2}D_{2}, the two bending modes occur at 511 and 538 cm^{−1}. The ν_{5}-ν_{4} difference bands are allowed and occur in the microwave, terahertz, and far-infrared wavelengths, with band origins at 117 cm^{−1} (3500 GHz) for ^{12}C_{2}H_{2} and 27 cm^{−1} (900 GHz) for ^{12}C_{2}D_{2}.

Two hundred and fifty-one ^{12}C_{2}D_{2} transitions, which are from ν_{5}-ν_{4}, (ν_{5}+ν_{4})-2ν_{4} and 2ν_{5}-(ν_{5}+ν_{4}) bands, have been measured in the 0.2-1.6 THz region, and 202 of them were observed for the first time. The precision of these measurements is estimated to be from 50 kHz to 100 kHz. A multistate analysis was carried out for the bending vibrational modes ν_{4} and ν_{5} of ^{12}C_{2}D_{2}, which includes the lines observed in this work and prior microwave, far-infrared and infrared data on the pure bending levels. Significantly improved molecular parameters were obtained for ^{12}C_{2}D_{2} by adding the new measurements to the old data set which had only 10 lines with microwave measurement precision. The experiments on ^{12}C_{2}H_{2} are in progress and ten P branch lines have been observed. We will present the ^{12}C_{2}H_{2} results to date.